OpenStudy (anonymous):

Which rule represents the sequence below? 13, 21, 29, 37, 45, ... Each term in the sequence is 10 less than the previous term. Each term in the sequence is 10 more than the previous term. Each term in the sequence is 8 less than the previous term. Each term in the sequence is 8 more than the previous term.

OpenStudy (anonymous):

Each term in the sequence is 8 more than the previous term. 13, 21, 29, 37, 45 13 21=13+8 29=21+8 37=29+8 45=37+8